OpenStudy (anonymous):

Find the exact coordinates of the vertex of the quadratic function g(x) = -2x^2+3x+1

OpenStudy (anonymous):

My solutions were .75 for the x coordinate and 4.375 for the y coordinate of the vertex. When I plug it into the calculator, the vertex is .42553191 for x and 2.6387505 and y

OpenStudy (anonymous):

@Kainui @iambatman @ganeshie8

OpenStudy (anonymous):

I used the vertex equation -b/2a and took that value and put it back into the original equation for find the y coord

OpenStudy (kainui):

The method seems right, except I'm confused. Why are you getting a second x-value? x=.75, that should be it if you're finding the vertex since there's only one.

OpenStudy (anonymous):

I got my first x value as .75 and my y value as 4.375

OpenStudy (anonymous):

I graphed the equation on the calculator and used the maximum to find the vertex and I get two completely different values

OpenStudy (kainui):

I just graphed it on wolfram alpha, you're right for your x-value but I think you might be getting your y-value wrong. http://www.wolframalpha.com/input/?i=vertex%20g(x)%20%3D%20-2x%5E2%2B3x%2B1&t=crmtb01

OpenStudy (anonymous):

-2(.75)^2 + 3(.75) + 1 = 1.125 + 2.25 + 1 = 4.375 that is my steps

OpenStudy (anonymous):

oh, I see what I did... I forgot to bring the negative over

OpenStudy (kainui):

Awesome, just allow yourself to take a little extra time with each problem on the test so that you don't miss a lot of tiny, simple things like that because you're too frantic.

OpenStudy (anonymous):

and on the calculator, I changed it from a -2x^2 to a 2x^2... I need to slow down when working through the problems

OpenStudy (anonymous):

I feel rushed a lot of the time because you only get 45 minutes on a test

OpenStudy (kainui):

How many questions is it?

OpenStudy (anonymous):

She usually does between 30 and 40 questions

OpenStudy (kainui):

I often times will come to a test, see the first question and freak out. Then I calmly forget about it and flip to the back of the test and worry about other problems I know that I can do. I don't know if that will help you or not, but then once I get comfortable with the test, then I'll return to the harder ones I skipped and I'll remember what to do, or I'll know that I can spend time trying to figure it out the longer way with whatever time I have left.

OpenStudy (anonymous):

Yeah, that is the same problem I have. This site has been great for helping me understand concepts and solving problems, but I freak out on the test and I still have not passed one of my tests but I feel I learn a lot.

OpenStudy (anonymous):

I mostly get freaked out by the time limit and then I try to work every problem and I end up skipping a few of them that would have got me more points.